The two wires in the segment below have the same diameter but different resistivities ?1 and ?2. Current I passes through the wire. If the ratio of resistivities is ?1/?2 = 4, what is the ratio of the electric fields in each wire, E1/E2?

P1
Answer #1

According to the given problem,

It's quite straightforward.

2 R=ρ

Since both segments have equal L and area, R11 = R22

==> R1/R2 = ρ12 = 4.

And I = V1/R1 = V2/R2,

where V1= E1L and V2 = E2L

==> V1/V2 = E1/E2 = R1/R2 = ρ12  = 4

==>  E1/E2 = 4